OpenStudy (anonymous):

Write a word problem that can be solved by working backward. Include at least two steps in the problem. Then, show the answer and how the problem can be solved by working backward.

OpenStudy (dan815):

If you add 2 to a number and divide by 3 and multiply by 2 you get 24 so what number did i start with? we can work backwards for this problem

OpenStudy (dan815):

last thing i did was multiply by 2 so divide by 2 24/2=12 then i divded by 3 to get here so i multiply by 3--> 12*3= 36 then i added 2 to this number so i subtract that 2, 36-2=34

OpenStudy (dan815):

so 34 was the answer i started with.. there working backwards lol

OpenStudy (dan815):

lets check to make sure this is right ((34+2)/3) *2 =24 yep works

OpenStudy (dan815):

can you write your own question similar to this now, maybe a little more creative too